Oscilloscope with RC circuits:

Explain with an example how you measure the frequency of a signal from oscilloscope by using (a) x-axis scale and (b) using y-axis scale?
(c) what is the difference between V(peak voltage) and Vrms (root-mean-square) of AC voltage source?
Answer #1

# measurement of frequency from oscilloscope

generally in oscilloscope shows Volt-time curve

here we have a oscilloscope curve -

X- axis show time and let time base is 2 ms/cm means wave takes 2 ms time to pass 2cm

   Y-axis show VOLTAGE GAIN let Y gain= 0.2 V/cm means per cm 2 volt gain is acquired

now for frequency measure 1st we find time period for this wave

time period = peak to peak at x-axis* time base

so in this example we have peak to peak value at x-axis is - 4cm

time period is= 4cm*2ms/cm = 8ms

now frequency is = 1/T = 1/0.008=125 Hz

c) Vpeak-to-peak  & Vrms

Vpeak-to-peak = it is define as the total voltage of waveform measure by the sum of maximum amplitude of voltage in -ve direction and +ve direction or -ve peak to +ve peak

Vrms=Root mean square is actually value of ac voltage at a instant of a time. mathematically it is define as the square root of mean of squared voltage

relation in both voltages is Vpeak to peak=Vrms×√2